OpenStudy (anonymous):

Help!! write in slope intercept form: slope = 1/2 through (-2,8) I know this far: y=1/2x .....

OpenStudy (anonymous):

This isn't a biology question and math isn't really my thing, but... I couldn't help but notice that the point you're given (-2, 8) is a set of two values: x and y. You need to find b in y = mx + b, and you already know that m is the slope (1/2 in this problem) so how do you think could could solve for b?

OpenStudy (anonymous):

oh oops, meant to put this on the math section

OpenStudy (anonymous):

Don't worry about it. Do you get what I was hinting at in my response, though?

OpenStudy (anonymous):

yes I know what you are saying but the problem is, where do i start graphing the line? i know how to graph lines using slope but i don't know what point to start on

OpenStudy (anonymous):

I'm saying that you can figure out the y intercept (b in y = mx + b) without graphing it. You have m, and you have x and y for a point on the line, so you can find b. Once you've got that, you can draw the graph, because b will tell you where to start on the y axis.

OpenStudy (anonymous):

do you know what the y-intercept is? i don't know what you mean.. well i do but i don't.

OpenStudy (anonymous):

The y-intercept is just the point where the line of the graph crosses the y axis. Like this: 

OpenStudy (anonymous):

i know this

OpenStudy (anonymous):

The equation of a line is y = mx + b Your two variables are x and y -- they change as you go along the line, though they stay in a fixed relationship defined by the equation. m is the slope b is the y-intercept In the question that you posted, you have m given as 1/2, and for an arbitrary point on the line, you have x = -2 and y = 8. Slot those into your equation y = mx + b and you can simply solve for b. Once you have that, the easiest way to draw the graph is just to connect the y-intercept that you calculated with the point (-2, 8) with a straight line.

OpenStudy (anonymous):

Not the right section but point slope form is simply\[y-y _{1}=m(x-x _{1})\] where y sub 1 and x sub 1 are the coordinates of the point given and m is your slope. so substitute 8 for y sub 1, (-2) for x sub 1, zero for x and 1/2 for m

OpenStudy (anonymous):

solving for y will yield your y-intercept, b for use in slope-intercept form

OpenStudy (anonymous):

@kickserve He/she needs slope-intercept form, though, which is y = mx + b

OpenStudy (anonymous):

@Calliope I realized that, however in order to solve for be without usage of a graph, using point slop as a stepping stone is very straightforward. Point slope is simply used to find b for y=mx+b

OpenStudy (anonymous):

Just seems a bit indirect when he/she can just plug y, m and x into y = mx + b and get b right away.

OpenStudy (anonymous):

very true, just illustrating another method as several professors often like this (notably more annoying) route

OpenStudy (anonymous):

Fair enough.